OpenStudy (anonymous):

The value of y varies directly with the value of x, and y = –21 when x = 3. Which equations represent the relationship between y and x? Choose exactly two answers that are correct.

OpenStudy (igreen):

Answers?

OpenStudy (anonymous):

-7x=y y/x=-7 x/y=7 xy=-7

OpenStudy (igreen):

y = kx This is the formula for direct variation. Plug in y = -21 and x = 3: -21 = k(3) -21 = 3k Divide 3 to both sides, what's -21 / 3?

OpenStudy (anonymous):

-7?

OpenStudy (igreen):

Yes, so -7 is our constant, plug that back in the equation: y = kx y = -7x This is one of our equations.

OpenStudy (igreen):

Now divide 'x' to both sides.

OpenStudy (anonymous):

-2.3?

OpenStudy (igreen):

No, it doesn't equal anything, we're just moving 'x' to the other side by dividing.

OpenStudy (anonymous):

Oh... Well up above you said x=3. Sorry D:

OpenStudy (igreen):

\(y = -7x\) \(\dfrac{y}{x} = \dfrac{-7x}{x}\) The x's on the right cancel out: \(\dfrac{y}{x} = \dfrac{-7 \cancel x}{\cancel x}\) Leaving us with: \(\dfrac{y}{x} = -7\)

OpenStudy (igreen):

Those are our 2 equations :)

OpenStudy (igreen):

\(y = -7x\) and \(\dfrac{y}{x} = -7\)

OpenStudy (anonymous):

Thank you! :D